OpenStudy (anonymous):

Suppose X is a normally distributed random variable with unknown mean μ and unknown standard deviation σ. Assume a random sample of X values, of size n = 25, is available. Let s^2 denote the sample variance. If σ^2=40, what is P(S^2≤23.08)?

OpenStudy (john_es):

\[P(S^2\leq23.08)=P(\frac{nS^2}{\sigma^2}\leq\frac{25\cdot23.08}{40})\] \[P(\chi^2_{24}\leq14.43)=0.6\] You must do an interpolation in order to get the value, or an online calculator.